A passenger train's speed is 60mi/h and a freight trains speed is 40mi/h. the passenger train travels the same distance in 1.5 hour less time then the freight train. How long does each train take to make the trip?
I know d= r*t and t=d/r, but no where in the problem does it say what the overall time of the trip was, just that one train was 1.5 hr less. Please help.

Answer by mananth(16946)

let distance be x
passenger train speed = 60 mi/h
time pasanger train = x/60
Similarly time freight train = x/40
Difference between the two timings is 1.5
x/40 -x/60 = 1.5
LCD = 120
3x-2x=120*1.5
x= 180 miles
Passenger train takes 180/60 = 3 hours
Freight tain takes 180/40 = 4.5 hours